# Conflict Adaptation in Vision-Language Models

### v1

@@ Abstract
A signature of human cognitive control is conflict adaptation: improved performance on a high-conflict trial following another high-conflict trial. This phenomenon offers an account for how cognitive control, a scarce resource, is recruited. Using a sequential Stroop task, we find that 12 of 13 vision-language models (VLMs) tested exhibit behavior consistent with conflict adaptation, with the lone exception likely reflecting a ceiling effect. To understand the representational basis of this behavior, we use sparse autoencoders (SAEs) to identify task-relevant “supernodes” in InternVL 3.5 4B. Partially overlapping supernodes emerge for text and color in both early and late layers, and their relative sizes mirror the automaticity asymmetry between reading and color naming in humans. We further isolate a conflict-modulated supernode in layers 24–25 whose ablation significantly increases Stroop errors while minimally affecting congruent trials.
Figure 1: Sequential Stroop task design. Left: an incongruent trial followed by another incongruent trial. Right: a congruent trial followed by an incongruent trial.
@@ 1 Introduction
A signature of human cognitive control is conflict adaptation: improved performance on incongruent trials when they follow other incongruent trials compared to congruent trials. This effect offers an account for how cognitive control, a scarce resource, is recruited botvinick2001conflict ; gratton1992optimizing . Whether VLMs exhibit conflict adaptation, and through what mechanisms, remains poorly understood. Understanding the emergence of such adaptation would inform theories of goal-directed processing across artificial and biological systems.
@@ 2.2 Task-Relevant Features
To understand how VLMs implement conflict processing, we leverage transcoders dunefsky2024transcoders , a variant of SAEs bricken2022monosemanticity ; cunningham2023sparse , which seek to decompose dense, polysemantic neural computations into sparse, monosemantic features that respond to specific patterns. We focus our analysis on InternVL 3.5 4B (pretrained), whose language model component, Qwen 3 4B yang2025qwen3 , has available transcoders circuit-tracer . While applying base language model transcoders to a VLM with modified weights arguably introduces potential concerns, our causal ablation experiments show that the extracted features remain functionally meaningful and interpretable.
To isolate features of interest, we construct “summary tensors” by averaging sparse activations across specific trial types and computing differences. We then apply coactivation-based grouping introduced by Deng et al. deng2025sparse to these summary tensors. This method constructs inter-layer feature networks based on temporal coactivation across tokens and extracts connected components, which we term “supernodes”, from the networks. We validate supernode importance through causal ablation: setting supernode features to zero during forward passes and measuring output distribution changes.
@@ 2.2.2 Conflict-Modulated Features
To identify conflict-related features, we compute summary tensor A(c1 != t1, c2 != t2) - A(c1 == t1, c2 != t2). Feature coactivation revealed a supernode in layers 24-25 showing more elevated activation under II than CI (Figure 6). Unlike the color and text features, which are sparse and have interpretable descriptions (top activating texts largely mention the corresponding colors), the conflict-modulated features are dense—activating frequently across diverse contexts—and lack clear interpretable descriptions. This suggests they serve general rather than task-specific functions.
@@
Ablating this supernode increases Stroop errors substantially: 3.38-fold on CI trials (17.5% to 59.2%) and 8.33-fold on II trials (2.5% to 20.8%), with minimal effect on congruent trials (Figure 7). Notably, post-ablation, the model frequently outputs the word “ink” instead of the second color, a failure mode entirely absent pre-ablation. Notably, post-ablation, the model sometimes repeatedly outputs the word “ink” in place of the second color, a failure mode entirely absent pre-ablation.
@@ 3 Limitations
Task comprehension presents a potential confound in cognitive evaluations of language models hu2024auxiliary ; hu-lewis-2025-language . Although we provide thorough instructions and all models demonstrate baseline accuracy, we cannot rule out that conflict may help models better understand task requirements. What appears as conflict adaptation may therefore partly reflect conflict clarifying what the task demands, rather than purely representing adaptive control mechanisms.

### v2

@@ Abstract
A signature of human cognitive control is conflict adaptation: improved performance on a high-conflict trial following another high-conflict trial. This phenomenon offers an account for how cognitive control, a scarce resource, is recruited. Using a sequential Stroop task, we find that 12 of 13 vision-language models (VLMs) tested exhibit behavior consistent with conflict adaptation, with the lone exception likely reflecting a ceiling effect. To understand the representational basis of this behavior, we use sparse autoencoders (SAEs) to identify task-relevant supernodes in InternVL 3.5 4B. Partially overlapping supernodes emerge for text and color in both early and late layers, and their relative sizes mirror the automaticity asymmetry between reading and color naming in humans. We further isolate a conflict-modulated supernode in layers 24–25 whose ablation significantly increases Stroop errors while minimally affecting congruent trials.
Figure 1: Sequential Stroop task design. Left: a congruent trial followed by an incongruent trial. Right: an incongruent trial followed by another incongruent trial.
@@ 1 Introduction
A signature of human cognitive control is conflict adaptation: improved performance on incongruent trials when they follow other incongruent trials compared to congruent trials. This effect offers an account for how cognitive control, a scarce resource, is recruited botvinick2001conflict; gratton1992optimizing. Whether VLMs exhibit conflict adaptation, and through what mechanisms, is poorly understood. Understanding the emergence of such adaptation would inform theories of goal-directed processing across artificial and natural minds.
@@ 2.2 Task-Relevant Features
To understand how VLMs implement conflict processing, we leverage transcoders dunefsky2024transcoders, a variant of SAEs bricken2022monosemanticity; cunningham2023sparse, which seek to decompose dense, polysemantic neural activations into sparse, monosemantic features that respond to specific patterns. We focus our analysis on InternVL 3.5 4B (pretrained), whose language model component, Qwen 3 4B yang2025qwen3, has available transcoders circuit-tracer. While applying base language model transcoders to a VLM with modified weights arguably introduces potential concerns, our causal ablation experiments show that the extracted features remain functionally meaningful and interpretable.
To isolate features of interest, we construct “summary tensors” by averaging sparse activations across specific trial types and computing differences. We then apply coactivation-based grouping introduced by Deng et al. deng2025sparse to these summary tensors. This method constructs inter-layer feature networks based on temporal coactivation across tokens and extracts connected components, which we term “supernodes”, from each network. We validate supernode importance through causal ablation: setting supernode features to zero during forward passes and measuring output distribution changes.
@@ 2.2.2 Conflict-Modulated Features
To identify conflict-related features, we compute summary tensor A(c1 != t1, c2 != t2) - A(c1 == t1, c2 != t2). Feature coactivation revealed a supernode in layers 24-25 showing more elevated activation under II than CI (Figure 6). Unlike color and text features, which are sparse and interpretable (top activating texts largely mention the corresponding colors), the conflict-modulated features are dense—activating frequently across diverse contexts—and lack clear interpretable descriptions. This suggests they serve general rather than task-specific functions.
@@
Ablating this supernode increases Stroop errors substantially: 3.38-fold on CI trials (17.5% to 59.2%) and 8.33-fold on II trials (2.5% to 20.8%), with minimal effect on congruent trials (Figure 7). Notably, post-ablation, the model sometimes repeatedly outputs the word “ink” in place of the second color, a failure mode entirely absent pre-ablation.
@@ 3 Limitations
Task comprehension presents a potential confound in cognitive evaluations of language models hu2024auxiliary; hu-lewis-2025-language. Although we provide thorough instructions and all models demonstrate baseline accuracy, we cannot rule out that conflict may help models better understand task requirements. What appears as conflict adaptation may therefore partly reflect conflict clarifying what the task demands.